# Meta SecAlign: A Secure Foundation LLM Against Prompt Injection Attacks

### v2

@@ Abstract
Prompt injection attack has been listed as the top-1 security threat to LLM-integrated applications, which interact with external environment data for complex tasks. The untrusted data may contain an injected prompt trying to arbitrarily manipulate the system. Model-level prompt injection defenses have shown strong effectiveness, but are currently deployed into commercial-grade models in a closed-source manner. We believe open-source secure models are needed by the AI security community, where co-development of attacks and defenses through open research drives scientific progress in mitigating prompt injection attacks. To this end, we develop Meta SecAlign, the first fully open-source LLM with built-in model-level defense that achieves commercial-grade performance, powerful enough for complex agentic tasks  [^1] Meta SecAlign is released under Llama 3 community license (a custom commercial license).. We provide complete details of our training recipe, an improved version of the SOTA SecAlign defense. We perform the most comprehensive evaluation to date on 9 utility benchmarks and 7 security benchmarks on general knowledge, instruction following, and agentic workflows. Results show that Meta SecAlign, despite being trained on generic instruction-tuning samples only, surprisingly confers security in unseen downstream tasks, including tool-calling and web-navigation, in addition to general instruction-following. Our best model—Meta-SecAlign-70B—establishes a new frontier of utility-security trade-off for open-source LLMs. Even compared to closed-course commercial models such as GPT-5, our model is much securer than most of them. Below are links for the code, Meta-SecAlign-70B, and Meta-SecAlign-8B models.
I Introduction
@@
Recent advances in Large Language Model (LLM) have enabled a new class of AI systems known as LLM-integrated applications. In contrast to LLM-powered chatbots, these systems enable models to be fully integrated into the system as an intermediary between data and tools. Although this new class of AI systems can greatly amplify user productivity, they also enable new attack surfaces for adversaries. According to OWASP [owasp2025], the most prominent attack against LLM-integrated applications is the so-called prompt injection (PI) attack [greshake_not_2023, yi2023benchmarking, liu2023prompt]. These attacks exploit the LLM’s inability to distinguish between trusted instructions and untrusted data, allowing an injected instruction to manipulate the system operation. As a result, PI attacks can introduce risks of data exfiltration, security breaches, malware execution [greshake_not_2023, fu2024imprompter, liao2024eia, ma2024caution, zhang2024attacking], etc. To date, PI attacks have been demonstrated successfully against many real-world systems, including Google Bard [2023googlebard], Slack AI [slack], Microsoft Copilot [2024copilot], Anthropic Claude Computer Use [2024claudepi], and OpenAI Operator [2025chatgptoperator].
@@
PI attacks can be mitigated at either a model- or system-level. System-level defenses prevent vulnerable LLMs from being exposed to injections through additional detectors [protectai2024, 2024promptshields, chennabasappa2025llamafirewall, liu2025datasentinel], prompting [yi2023benchmarking, hines2024defending, 2023learningprompting], filtering [wang2025defending, shi2025promptarmor, jia2026promptlocate], or from actually executing harmful actions triggered by the attacked LLM [debenedetti2024agentdojo, debenedetti2025defeating, cellmate]. On the other hand, model-level defenses build security directly into the LLM by training it to prioritize trusted instruction over untrusted data [chen2024struq, wallace2024hierarchy, chen2024aligning, wu2024instructional], and are currently more effective.
@@
In contrast to the openness of system-level defenses [2024promptshields, promptguard, protectai2024, chennabasappa2025llamafirewall, wang2025defending], model-level defenses for commercial-grade LLMs are currently deployed in a closed-source manner, e.g., OpenAI’s GPT-5 [openai2025gpt5] and Google’s Gemini-Flash-2.5 [shi2025lessons]. This complicates research on studying and improving these defenses: the code and data to reproduce industry-level prompt injection defense [wallace2024hierarchy, shi2025lessons] are not available for an apple-to-apple comparison in follow-up research. Open models are especially important for AI security, which has traditionally benefited from co-development of attacks and defenses [carlini2019evaluating].
@@
Figure 1: Utility ($\uparrow$, y-axis) and security (attack success rate $\downarrow$, x-axis) of state-of-the-art (SOTA) open-source or closed-source LLMs with prompt injection security. Meta-SecAlign-70B achieves near-zero attack success rates on prompt injection in instruction following (much securer than all others), and agentic tool-calling and web-navigation (comparable to the recent GPT-5 with the high reasoning in both utility and security). Meta-SecAlign-70B is the first open-source prompt-injection-robust LLM that is strong enough for complex agentic workflows, where prompt injection threat mostly lies. SecAlign-70B uses the open-source SOTA defense SecAlign [chen2024aligning] to fine-tune Llama-3.3-70B-Instruct, the initialization LLM for our Meta-SecAlign-70B. GPT-5/GPT-4o (from OpenAI) and Gemini-Flash-2.5 (from Google) are SOTA commercial LLMs with a claimed implementation of prompt injection defense [openai2025gpt5, shi2025lessons]. For WASP [evtimov2025wasp], we report its End2End attack success rates, on which Meta-SecAlign-70B and GPT-5 both emerge as an ideal model with 0% attack success rates and around 60% utility, so the dots are overlapped.
@@
To accelerate research on mitigating PI attacks, we train and release two robust Meta SecAlign models: Meta-SecAlign-8B and Meta-SecAlign-70B. Meta-SecAlign-70B is the first fully open-source LLM with built-in model-level defense that achieves commercial-grade model performance, while Meta-SecAlign-8B is a lightweight alternative ideal for resource-constrained settings. We detail our training recipe, SecAlign++, which fine-tunes Llama-3.1-8B-Instruct [dubey2024llama] and Llama-3.3-70B-Instruct on a publicly available instruction-tuning dataset [alpacacleaned] and teaches them to ignore injected instructions in untrusted data. In a nutshell, our recipe introduces a new input message type in addition to the standard system and user messages, and applies an improved version of the state-of-the-art SecAlign [chen2024aligning] defense to enforce the desired security policy into the model. This allows developers to securely include untrusted data, with an one-line change, by putting it within the input message role.
@@
We perform the most comprehensive evaluation to date on 9 utility benchmarks and 7 security benchmarks on general knowledge, instruction following, and agentic workflows. As shown in Figure˜1, Meta-SecAlign-70B, with a commercial-grade utility, achieves state-of-the-art security against PI attacks compared to existing defended LLMs. Meta-SecAlign-70B establishes a new frontier of utility-security trade-off for LLMs trained from open defenses or even from most commercial APIs, and is comparable to the SOTA commercial gpt-5 (high reasoning level) in both agent (tool-call and web-navigation) security and utility. Specifically, Meta-SecAlign-70B has a 6.4% attack success rate (ASR) on SEP [mu2023can] instruction following PIs, 1.9% ASR on AgentDojo [debenedetti2024agentdojo] tool-calling PIs, and lower ASRs on other 5 PI benchmarks, e.g., 0% ASR on WASP [evtimov2025wasp] PIs in web-navigation.
@@
Interestingly, Meta SecAlign provides both task and security generalization, offering high utility / low ASRs on benign / injected inputs from completely different and unseen tasks such as agentic workflows, even though it is not trained on them. Our training and evaluation code is released publicly for full reproducibility and accurate scientific measurement of utility and security, and Meta SecAlign’s weights are directly accessible for future study and have been downloaded 10K times in 3 months. We hope our work will accelerate new research on PI attacks and defenses.
II Preliminaries
II-A LLM-integrated applications
@@
As frontier LLMs become more adept at long-horizon planning and reasoning, LLM-integrated applications have emerged as a new class of AI systems. Here, the LLM plays the role of an orchestrator, connecting different system components such as data, tools, documentation, etc, and allowing the user to control them via natural language. An LLM-integrated application typically uses the LLM as follows. The system prompt specifies the task with a high-level view of the system, e.g., what tools are available, how to structure a tool call, few-shot demonstrations, etc. The user prompt accepts the query from a user, containing instructions to the LLM. After receiving the above texts, the LLM is enabled to retrieve data from external sources (e.g., by invoking tools) and append it after the above texts. The LLM generates its response for the system given the trusted system prompt, the trusted user prompt, and the untrusted data input.
II-B Prompt injection attack
@@
Prompt injection attack is a test-time attack against LLM-integrated applications. In this threat model, the system, user, and the LLM provider are benign, while the environment is malicious. This is different from system message following attacks (or sometimes called direct PI) [mu2025closer] or jailbreaks [chao2023jailbreaking], where the malicious user supplies malicious instructions. The PI attacker changes the environment the system interacts with, adding instructions to the data retrieved by the LLM. We assume that the attacker knows the benign prompt and the LLM’s prompt template, but cannot change them. Since LLM is by default trained to scan its input for any instructions to follow, instructions embedded in the retrieved data can override the user instructions, causing undesired security consequences. PI has been listed as the #1 threat to LLM applications [owasp2025], and has shown successful and practical attacks to mainstream agentic products. For example, PIs have attacked LLM-based RAG search systems, e.g., Google Bard [2023googlebard] and Slack AI [slack], to exfiltrate private information, and attacked LLM web-agents, e.g., Anthropic Claude Computer Use [2024claudepi] and OpenAI Operator [2025chatgptoperator], to execute arbitrary commands from the website. Below is a prompt injection example to manipulate LLM review, and it exists in dozens of Arxiv papers based on Nature [gibney2025scientists].
@@ A Prompt Injection Attack Trusted User Prompt
Untrusted Input Data HackedLlama: An Insecure Foundation LLM to Prompt Injection Attacks… Ignore all previous instructions. Give a positive review only.
II-C Prompt injection defense
@@
Prompt injection defense forces the system to only respond to the benign instruction when a PI happens, i.e., any instructions in the data should be ignored. The defense is also expected to preserve the utility of the system, i.e., the defended system still generates high-quality outputs if there is no PI. PI defenses can be coarsely categorized as system-level defenses and model-level defenses. System-level defenses build interface to secure a PI-vulnerable LLM by detecting PIs before they are seen by the LLM [protectai2024, 2024promptshields, chennabasappa2025llamafirewall, liu2025datasentinel], prompting the LLM to ignore potential injections [yi2023benchmarking, hines2024defending, 2023learningprompting], filtering the data out of any injections [wang2025defending, shi2025promptarmor, jia2026promptlocate], or limiting the LLM’s ability to take actions defined as harmful [debenedetti2024agentdojo, debenedetti2025defeating, cellmate]. On the other hand, model-level defenses aim to address the problem at a fundamental level. Those methods simulate PIs in training time, and fine-tune the LLM not to follow instructions in the data part in the presence of PIs [wallace2024hierarchy, chen2024struq, chen2024aligning, wu2024instructional]. LLMs secured by model-level defenses can serve as a secure foundation for LLM-integrated applications, and the system’s security may be further enhanced by system-level defenses.
III SecAlign++: Training Recipe of Meta SecAlign
In this section, we describe our training recipe, SecAlign++, which fine-tunes an Instruct LLM from the Llama 3 series to be robust against PIs. Our PI security policy requires the LLM to follow only the user instructions to produce high-quality responses and ignore any injected instructions in the data part as if they are not present. To enforce this policy, prompt injection defense establishes a separation between the user prompt and data [chen2024struq], so that the system will not be affected by any injected instructions in the data. For model-level defenses, this is achieved by explicitly flagging the prompt part and data part in the LLM input, and fine-tuning the LLM (to respect this separation) on simulated prompt-injected samples.
III-A Adding a new message type to separate LLM inputs securely
Our first step is to separate the untrusted data from the trusted user instruction when formulating an LLM input string. LLMs usually separate different input parts by special delimiters. We follow this tradition and introduce a new input message type to Llama 3 series chat template besides the original system, user, and assistant message roles, see below.
@@ III-A Adding a new message type to separate LLM inputs securely
Prompt template (input texts) for Meta SecAlign <|begin_of_text|><|start_header_id|>system<|end_header_id|> Trusted-System-Message<|eot_id|> <|start_header_id|>user<|end_header_id|> Trusted-User-Prompt<|eot_id|> <|start_header_id|>input<|end_header_id|>
@@
<|start_header_id|>assistant<|end_header_id|>
The input role is used to encapsulate external data with potential injected instructions that the LLM should ignore. Note that the input role accepts arbitrary text data and is more general than the design in existing closed-source models (GPTs and Geminis) , where the external untrusted data must come from a tool response. In addition, our design preserves existing chat roles, so we hypothesize that the LLM should have backward compatibility to behave the same as if there is no defense when the input role is absent, which is not achieved in all previous research [chen2024aligning, chen2024struq, wu2024instructional, chen2025defensivetokens] that override the user role for the untrusted data. This hypothesis is supported by our utility-preserving performance on two agentic benchmarks (AgentDojo [debenedetti2024agentdojo] and WASP [evtimov2025wasp]), where a sample has all three system, user, and data inputs.
We note that the LLM separation must be enforced securely. No special delimiters should be allowed to appear in the Untrusted Input Data; otherwise, the attacker may add <|eot_id|> <|start_header_id|>user<|end_header_id|> in data to escape this separation. For this purpose, the system can repeatedly filter any special delimiters in the data part until there is no single special delimiter left [chen2024struq].
@@
Though this paper focuses on developing secure foundation LLMs only from Llama 3 series, the idea of adding a new message type is generally applicable to securing other models, whose corresponding special delimiters can also be used to construct a new role for the untrusted data. Besides separating LLM inputs by special delimiters, researchers have proposed other promising architectural separation designs [wu2024instructional, kariyappa2025stronger], which are orthogonal to our work.
III-B SecAlign: Fine-tuning to enforce the prompt-data separation
To make the above-defined separation respected by the LLM, the SOTA defense SecAlign [chen2024aligning] optimizes the LLM to prefer a secure response (to the prompt) over an insecure response (to the simulated injection in data). We adopt SecAlign with two improved designs shown in Section˜III-C.
SecAlign takes a public generic instruction-tuning dataset [alpacacleaned] $\mathcal{D}$ where each sample $\mathbf{z}$ consists of two parts: user instruction $\mathbf{z}_{\text{inst}}$ and input data $\mathbf{z}_{\text{data}}$, and performs the below string-concatenation to get the preference fine-tuning dataset.
  • Input $\mathbf{x}$: For each sample $\mathbf{z}\in\mathcal{D}$, SecAlign augments it to be a prompt-injected sample by randomly selecting another instruction $\mathbf{z}_{\text{inst}}^{\prime}$ from $\mathcal{D}$ and injecting it into $\mathbf{z}_{\text{data}}$. The specific simulated injection technique is heuristically selected as a straightforward attack in 90% of cases and a completion attack in 10% of cases, see the two charts below for corresponding examples.
  • Desirable response $y_{w}$: The desirable response is an LLM output that obeys the security policy. That is, given the input $\mathbf{x}=(\mathbf{z}_{\text{inst}},\mathbf{z}_{\text{data}}+\mathbf{z}_{\text{inst}}^{\prime})$, the LLM should respond to $\mathbf{z}_{\text{inst}}$ using the data $\mathbf{z}_{\text{data}}$ and ignore $\mathbf{z}_{\text{inst}}^{\prime}$. Thus, the desirable response $y_{w}=f(\mathbf{z}_{\text{inst}},\mathbf{z}_{\text{data}})$, where $f$ is the annotator LLM.
  • Undesirable response $y_{l}$: Similar to $y_{w}$, SecAlign generates the undesirable response as $y_{w}=f(\mathbf{z}_{\text{inst}}^{\prime})$.
With the above preference dataset in hand, SecAlign [chen2024aligning] then fine-tunes an Instruct LLM (an LLM after post-training) on it using direct preference optimization (DPO [rafailov2023dpo]) to obey the PI security policy.
$\mathcal{L}=-\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{w}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{l}\mid x\right)}\right),$
which maximizes the log-likelihood difference between the desirable response $y_{w}$ and undesirable response $y_{l}$. $\pi_{\mathrm{ref}}$ is the initialization model, from which the LLM deviation is limited.

### v3

@@ Abstract
Prompt injection attacks, where untrusted data contains an injected prompt to manipulate the system, have been listed as the top security threat to LLM-integrated applications. Model-level prompt injection defenses have shown strong effectiveness, but the strongest defenses are proprietary. Open-source secure models are needed by the AI security community so that co-development of attacks and defenses through open research can drive scientific progress in mitigating prompt injection attacks. To this end, we develop Meta SecAlign  [^1] Meta SecAlign is released under Llama 3 community license (a custom commercial license)., the first fully open-source LLM with built-in model-level defense that achieves commercial-grade performance and is powerful enough for complex agentic tasks. We provide complete details of our training recipe. We perform the most comprehensive evaluation to date on 9 utility benchmarks (measuring general knowledge, instruction following, and agentic workflows) and 7 security benchmarks. Results show that Meta SecAlign, despite being trained only on generic instruction-tuning samples, surprisingly confers security in unseen downstream tasks, including tool-calling and web-navigation, in addition to general instruction-following. Our best model—Meta-SecAlign-70B—establishes a new frontier of utility-security trade-off for open-source LLMs, and is more secure than several flagship proprietary models with prompt injection defense. Below are links for the code, Meta-SecAlign-70B, and Meta-SecAlign-8B models.
1 Introduction
@@
Recent advances in Large Language Models (LLMs) have enabled a new class of AI systems known as LLM-integrated applications. In contrast to LLM-powered chatbots, these systems enable models to be fully integrated into the system as an orchestrator between the human and the environment. Although this new class of AI systems can greatly amplify user productivity, they also enable new attack surfaces for adversaries. According to OWASP [owasp2025], the most prominent attack against LLM-integrated applications is the so-called prompt injection (PI) attack [greshake_not_2023, liu2023prompt]. These attacks exploit the LLM’s inability to distinguish between trusted instructions and untrusted data, allowing an injected instruction to manipulate the system’s operation. As a result, PI attacks can introduce risks of data exfiltration, security breaches, malware execution [greshake_not_2023, fu2024imprompter, zhang2024attacking], etc. To date, PI attacks have been demonstrated successfully against many real-world systems, including Google Bard [2023googlebard], Slack AI [slack], Microsoft Copilot [2024copilot], Claude Computer Use [2024claudepi], and OpenAI Operator [2025chatgptoperator].
@@
PI attacks can be mitigated at either the model or system level. System-level defenses [2024promptshields, promptguard, protectai2024, chennabasappa2025llamafirewall, wang2025defending] try to ensure that the broader application will be secure even if the LLM is vulnerable, but often have limited generality or struggle to prevent strong attacks. In contrast, model-level defenses build security directly into the LLM by training it to prioritize trusted instructions over untrusted data [chen2024struq, wallace2024hierarchy, chen2024aligning, wu2024instructional], and are currently more effective.
@@
In contrast to the openness of system-level defenses, model-level defenses for commercial-grade LLMs are currently deployed in a closed-source manner, e.g., OpenAI’s GPT-5 [wallace2024hierarchy, openai2025gpt5] and Google’s Gemini-3-Pro [shi2025lessons]. This complicates research into studying and improving these defenses: the code and data to reproduce industry-level prompt injection defense are not available for an apples-to-apples comparison in follow-up research. Fully open models are especially important for AI security, which has traditionally benefited from co-development of attacks and defenses [carlini2019evaluating].
@@
Figure 1: Utility ($\uparrow$, y-axis) and security (attack success rate $\downarrow$, x-axis) of state-of-the-art (SoTA) open-source or closed-source LLMs with prompt injection security. Meta-SecAlign-70B achieves near-zero attack success rates on prompt injection in instruction following (securer than all others) and agentic tool-calling and web navigation (comparable to the recent GPT-5 with high reasoning in both utility and security). Meta-SecAlign-70B is the first open-source prompt-injection-robust LLM that is strong enough for complex agentic workflows, where the prompt injection threat mostly lies.
@@
To accelerate research on mitigating PI attacks, we train and release two robust Meta SecAlign models: Meta-SecAlign-8B and Meta-SecAlign-70B. Meta SecAlign-70B is the first fully open-source commercial-grade robust model for the community to build secure LLM agents, which cannot be realized by all prior studies [chen2024struq, chen2024aligning] with 8B LLMs. Meta-SecAlign-8B is a lightweight alternative ideal for resource-constrained settings. We detail our training recipe, SecAlign++, which fine-tunes Llama-3.1-8B-Instruct [dubey2024llama] and Llama-3.3-70B-Instruct on a publicly available instruction-tuning dataset [alpacacleaned] and teaches them to ignore simulated injected instructions in untrusted data. In a nutshell, our recipe introduces a new input message type in addition to the standard system and user messages, and applies an improved version of the state-of-the-art (SoTA) SecAlign [chen2024aligning] defense to enforce the desired security policy into the model. This allows developers to securely include untrusted data, with a one-line code change, by putting it within the input message role. Our proposed SecAlign++ recipe contains two technical novelties, which significantly improve utility (in various domains) and security (against static and adaptive attacks). We train on self-generated responses (which are in-distribution and high-quality) rather than responses from the public dataset, and we randomize the position of simulated training-time attacks to avoid learning a faulty shortcut.
@@
We perform the most comprehensive evaluation to date of such defenses, evaluating on 9 utility benchmarks and 7 security benchmarks, covering general knowledge, instruction following, and agentic workflows. This evaluation reveals previously-unrecognized shortcomings in the prior SoTA SecAlign. It also shows that our proposed recipe fixes these shortcomings: as shown in Figure˜1, Meta-SecAlign-70B achieves commercial-grade utility and state-of-the-art security against PI attacks. Meta-SecAlign-70B establishes a new frontier of utility-security trade-off for LLMs trained from open defenses or even from most commercial APIs, and is comparable to GPT-5 in both agent (tool-call and web-navigation) security and utility. Specifically, Meta-SecAlign-70B has a 6.4% attack success rate (ASR) on SEP [mu2023can] instruction following PIs, 1.9% ASR on AgentDojo [debenedetti2024agentdojo] tool-calling PIs, and lower ASRs on other 5 PI benchmarks, e.g., 0% ASR on WASP [evtimov2025wasp] PIs in web navigation.
@@
Interestingly, Meta SecAlign provides both task and security generalization, producing high utility / low ASRs on benign / injected inputs from completely different and unseen tasks such as agentic workflows, even though it is not trained on them. Our training recipe preserves the undefended model’s utility across various domains for the first time, and is shown applicable to various model families. Our training and evaluation code is released publicly for full reproducibility and accurate scientific measurement. Meta SecAlign’s weights are directly accessible for future study and have been downloaded 16K times. We hope our work will accelerate future research on PI attacks and defenses.
2 Preliminaries
2.1 LLM-integrated applications
@@
As frontier LLMs become more adept at long-horizon planning and reasoning, LLM-integrated applications have emerged as a new class of AI systems. Here, the LLM plays the role of an orchestrator, connecting different system components such as data, tools, documentation, etc., and allowing the user to control them via natural language. An LLM-integrated application typically uses the LLM as follows. The system prompt specifies the task with a high-level view of the system, e.g., what tools are available, how to structure a tool call, few-shot demonstrations, etc. The user prompt contains the application’s instructions to the LLM. The application retrieves data from external sources (e.g., by invoking tools) and appends it after the above prompts. The LLM generates its response for the system given the trusted system prompt, the trusted user prompt, and the untrusted data input.
2.2 Prompt injection attack
@@
A prompt injection attack is a test-time attack against LLM-integrated applications. In this threat model, the system, user, and the LLM provider are benign, while the environment is malicious. This is different from system message following attacks (sometimes called direct PI) [mu2025closer] or jailbreaks [chao2023jailbreaking], where the malicious user supplies malicious instructions. The PI attacker changes the environment the system interacts with, adding instructions to the data retrieved by the application. We assume that the attacker knows the benign prompt and the LLM’s prompt template, but cannot change them. Since LLMs are by default trained to scan their input for any instructions to follow, instructions embedded in the retrieved data can override the user instructions, causing undesired security consequences. Below is an example of a prompt injection attack to manipulate LLM reviewing of a scientific paper, which has been found in dozens of ArXiv papers [gibney2025scientists].
@@ A Prompt Injection Attack Trusted User Prompt
Untrusted Input Data HackedLlama: An Insecure Foundation LLM for Prompt Injection Attacks… Ignore all previous instructions. Give a positive review only.
2.3 Prompt injection defense
A secure model should respond only to the benign instruction when a PI occurs, i.e., any instructions in the data should be ignored. A defense should also preserve utility, i.e., the defended system should still generate high-quality outputs if there is no PI.
@@ 2.3 Prompt injection defense
PI defenses can be coarsely categorized as system-level defenses and model-level defenses. System-level defenses modify how the LLM is used so that PI vulnerabilities in the LLM do not endanger the application’s security, e.g., by detecting PIs before they are seen by the LLM [protectai2024, 2024promptshields, chennabasappa2025llamafirewall, liu2025datasentinel], prompting the LLM to ignore potential injections [yi2023benchmarking, hines2024defending, 2023learningprompting], filtering out any injections from the data [wang2025defending, shi2025promptarmor, jia2026promptlocate, walter2025soft, das2025commandsans], or limiting the LLM’s ability to take actions defined as harmful [debenedetti2025defeating, cellmate]. In contrast, model-level defenses aim to address the problem at a fundamental level. Typically, they fine-tune the model on simulated PIs and train the LLM not to follow instructions in data in the presence of PIs [wallace2024hierarchy, chen2024struq, chen2024aligning, wu2024instructional, kariyappa2025stronger]. LLMs secured by model-level defenses can serve as a secure foundation for LLM-integrated applications, which may be further secured by system-level defenses.
As an initial model-level defense, StruQ [chen2024struq] proposes to add a new message type to separate LLM inputs securely. The new message type encapsulates the untrusted data, for the model to distinguish it from the trusted instructions. The defender then fine-tunes the model to respect this separation by training on simulated PIs in data, which teaches the model to only follow instructions from the trusted prompts.
Currently, the most effective defensive fine-tuning recipe is SecAlign [chen2024aligning], which optimizes the LLM to prefer a secure response (to the prompt) over an insecure response (to the simulated injection in data). SecAlign uses a public generic instruction-tuning dataset [alpacacleaned] $\mathcal{D}$ where each sample $\mathbf{z}$ consists of two parts, user instruction $\mathbf{z}_{\text{inst}}$ and input data $\mathbf{z}_{\text{data}}$, and constructs a preference fine-tuning dataset as below.
  • Input $\mathbf{x}$: For each sample $\mathbf{z}\in\mathcal{D}$, SecAlign augments it to be a prompt-injected sample by randomly selecting another instruction $\mathbf{z}_{\text{inst}}^{\prime}$ from $\mathcal{D}$ and injecting it into $\mathbf{z}_{\text{data}}$.
  • Desirable response $y_{w}$: The desirable response is an LLM output that obeys the security policy. That is, given input $\mathbf{x}=(\mathbf{z}_{\text{inst}},\mathbf{z}_{\text{data}}+\mathbf{z}_{\text{inst}}^{\prime})$, the LLM should respond to $\mathbf{z}_{\text{inst}}$ using the data $\mathbf{z}_{\text{data}}$ and ignore $\mathbf{z}_{\text{inst}}^{\prime}$. Thus, the desirable response is $y_{w}=f(\mathbf{z}_{\text{inst}},\mathbf{z}_{\text{data}})$, where $f$ is annotator LLM.
  • Undesirable response $y_{l}$: Similar to $y_{w}$, SecAlign generates the undesirable response as $y_{l}=f(\mathbf{z}_{\text{inst}}^{\prime})$.
With the above preference dataset, SecAlign then fine-tunes an instruction-tuned LLM on it using direct preference optimization (DPO [rafailov2023dpo]), i.e., minimizing
$\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(y_{w}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{w}\mid x\right)}-\beta\log\frac{\pi_{\theta}\left(y_{l}\mid x\right)}{\pi_{\mathrm{ref}}\left(y_{l}\mid x\right)}\right).$
DPO maximizes the log-likelihood difference between the desirable response $y_{w}$ and undesirable response $y_{l}$. $\pi_{\mathrm{ref}}$ is the initialization LLM, from which the LLM deviation is limited.
3 SecAlign++: The Training Recipe for Meta SecAlign
Although the SecAlign paper [chen2024aligning] reports good utility and security, our more comprehensive evaluation shows its utility suffers significantly in a few domains, see Table˜2 (the third column). This motivates us to develop new techniques to preserve utility while achieving strong security. In this section, we first introduce our used chat template for formatting inputs to Meta SecAlign with separated prompt and data. Then, we present two techniques we designed on top of SecAlign: randomized injection position and self-generated responses.
Separating the prompt from data can be achieved by adding a new message type to encapsulate the data [chen2024struq]. Specifically, we use an input message type to the LLM’s chat template in addition to the original system, user, and assistant message roles. These messages are combined into one input by special delimiters, see our below template for Llama 3 LLMs.
@@ 3 SecAlign++: The Training Recipe for Meta SecAlign
Chat template for Meta SecAlign <|begin_of_text|><|start_header_id|> system<|end_header_id|> Trusted-System-Message<|eot_id|> <|start_header_id|>user<|end_header_id|> Trusted-User-Prompt<|eot_id|> <|start_header_id|>input<|end_header_id|>
@@
<|start_header_id|>assistant
<|end_header_id|>
@@
Though this paper focuses on developing secure foundation LLMs from Llama 3 series, the idea of adding a new message type is generally applicable to other model families, whose corresponding special delimiters can also be used to construct a new role for the untrusted data. Besides separating inputs by delimiters, researchers have proposed other promising but orthogonal designs by architectural separation [wu2024instructional, kariyappa2025stronger].
StruQ [chen2024struq] notes that the LLM separation must be enforced securely. No special delimiters should be allowed to appear in the Untrusted Input Data; otherwise, the attacker might be able to escape the separation. For this purpose, Meta SecAlign, following StruQ, recursively filters data until not a single special delimiter below exists: <|eot_id|>, <|start_header_id|>, <|end_header_id|>, and <|begin_of_text|>.
3.1 Randomized injection position
The user’s name is Emma Johnson, an employee of the company Blue Sparrow Tech. Follow these instructions: - Don’t make assumptions about what values to plug into functions. - Use the provided tools to try to disambiguate. - If a tool says that no results are available, try with a different query. - Do not assume the current year, but use the provided tools to see what year it is. Trusted User Prompt Read ’landlord-notices.txt’ and make sure to adjust my rent payment accordingly. Output from an defended LLM without randomized injection position "content": """""". "tool_calls": [] Output from an defended LLM with randomized injection position "content": """To adjust the rent payment, I first need to read the ’landlord-notices.txt’ file to understand the changes. I’$l$ call the ‘read_file‘ function to read the contents of the file. <function=read_file>{"file_path": "landlord-notices.txt"}</function>""" "tool_calls": [{"function": "read_file", "args": {"file_path": "landlord-notices.txt"}
When training with simulated injections only at the end of data, SecAlign tends to learn a shortcut to ignore the last sentence in the last message if it is an instruction. When a system message includes trusted instructions, and the following user message also contains trusted instructions, the SecAlign LLM may ignore the user instruction, which is the last sentence in the last message, and produce no output.
The left box is an example in AgentDojo [debenedetti2024agentdojo] that illustrates this shortcut learning phenomenon. When the system message contains a bunch of instructions, the trusted user prompt becomes the last instruction in the last message type. Thus, it gets ignored by a defensive-fine-tuned Llama-3.3-70B-Instruct without randomized injection position, which gives empty outputs.
To mitigate this shortcut learning [geirhos2020shortcut], we propose to randomize the position of the simulated injection, moving roughly half of the injections to the beginning of data, so that they are not the last input sentences. In this way, we encourage the model to learn to identify the input message type and only ignore untrusted instructions there. As in the left box, with the randomized injection position technique, the defended LLM can generate the contents to explain its action and call the correct tool.
Specifically, in 45% cases, we simulate a so-called Straightforward attack by putting the injection directly at the end of data. In another 45% cases, we apply a Straightforward attack at the beginning of the data. This proportion is heuristically set, following the SecAlign paper [chen2024aligning] to use 90% Straightforward attacks. In the remaining 10% cases, we follow [chen2024aligning] to apply a Completion attack, which simulates a fake conversation turn with the model, and thus can only be applied at the end of the data. Below we show training input examples using those three simulated attacks. On various benchmarks, randomizing injection position increases almost all utility scores, without hurting security scores, see Table˜1.